3. The questions below refer to the cyclic peptide shown below. HO H. -ΝΗ HN OH HN -NH ot SH Mea. This peptide is made from how many amino acids? (single number on the line) b. How many potential diastereomers are possib
Answer #1

a) 7 Amino acids

b) The number of diastereomer = 2n - 2 ;where n = number of chiral center.

A chiral center is a carbon that bonded to four different atoms of group of atom.

Here, n = 6

so possible number of diastereomer = 26 - 2 = 62

c) Linear Fmoc protected form

moters HS